OpenStudy (anonymous):

Please help me How do I write Y-2=4(x+1) into point slope form And 2x+5y=15 Please help me and thank you 

OpenStudy (anonymous):

I just don't get how to plug the numbers into the formula

OpenStudy (anonymous):

\[y-2=4(x+1) \] distribute the \(4\) \[y-2=4x+4\] then add \(2\) \[y=4x+6\]

OpenStudy (anonymous):

it is not a matter of plugging in to a formula there is no formula it is "solve for \(y\)"

OpenStudy (anonymous):

But the question says write in point slope form

OpenStudy (anonymous):

yes, what that means is "solve for \(y\)" to make it look like \[y=mx+b\] but it is not plugging in to a formula, it is using algebra to get \(y\) by itself on one side of the equal sign

OpenStudy (anonymous):

Oh ok thanks 😊 It's just algebra is very confusing to me

OpenStudy (anonymous):

How would I solve 2x+5y=15 into slope intercept form